Now we introduce the role of income taxes into the Keynesian system. U.S. taxes consist of autonomous taxes and income taxes. Autonomous taxes such as sales taxes and property taxes do not depend on income, whereas income taxes such as personal income taxes and corporate income taxes depend on the level of income. When we introduce income taxes into the model, the multipliers are modified as follows:

Assuming that the U.S. MPC is 0.9, and the average income tax rate is 0.3 (30%), calculate the multipliers

The investment multiplier = 1/[1 - MPC(1 - Tax rate)].
Answer #1

The formula for investment multiplier uses the average tax rate which is 0.3 and at the marginal propensity to consume is 0.9

Multiplier = 1/(1-0.9*(1-0.3))

= 1/(1-0.63)

= 1/0.37

= 2.703

This is the investment multiplier